OpenStudy (anonymous):

Solve the equation: x2 + 8x + 16 = 49 Enter the smallest zero value.

OpenStudy (nenadmatematika):

x^2+8x-33=0 (x+11)(x-3)=0 x=-11 or x=3 so 3 is the answer

OpenStudy (nenadmatematika):

sorry -11 is the smallest :D

OpenStudy (anonymous):

thank you guys! <3

OpenStudy (nenadmatematika):

you're welcome :D